OpenStudy (anonymous):

Kermit makes $6.0 per hour as a lifeguard and $10.0 per hour as a delivery guy. One week he worked a total of 20 hr and made $80.0 how many hours did he spend as a lifeguard

OpenStudy (anonymous):

ok, ill be right back sorry!

OpenStudy (anonymous):

I think you made a mistake...If Kermit spend 100% of his/her time (20 hours) on the lower job ($6), he/she would make $120 (The lowest possible salary for that week)....

OpenStudy (anonymous):

ya, so this is impossible to solve with the time you gave

OpenStudy (anonymous):

Yeah I made a mistake it sould have read: Kermit makes $6.0 per hour as a lifeguard and $10.0 per hour as a delivery guy. One week he worked a total of 20 hrs and made $180.0 how many hours did he spend as a lifeguard

OpenStudy (anonymous):

okay, thats better, no wonder

OpenStudy (anonymous):

my mom typed that so..hahaha shes blond

OpenStudy (anonymous):

5 hours as a life guard

OpenStudy (anonymous):

15 hours as a delivery guy

OpenStudy (anonymous):

5*6=30 15*10= 150 150+30=180 *